# Pangenome-guided sequence assembly via binary optimisation

### v1

@@ 4.7 Mapping QUBO problems to quantum optimisation
The circuit is constructed from a layer of Hadamard gates followed by alternating cost and mixer layers. The cost layer is $U_{C}(\gamma)=\exp(-i\gamma H)$. There are many choices for a suitable mixer layer; we use $U_{M}(\beta)=\exp(-i\beta\sum_{i}X_{i})$. The number of such layers is a hyperparameter $p$. In the limit as $p\rightarrow\infty$, QAOA approaches adiabatic quantum computing, which is known to be universal for quantum computing [28]. In contrast, for $p=1$, QAOA is known to be classically simulable under certain circumstances [29]. In the absence of noise, increasing $p$ should increase the convergence rate of QAOA. In practice, the effects of noise mean that a moderate $p$ value of between 2 and 20 layers is most likely to be effective.
@@ 4.8 Classical simulation of QAOA
We find that the convergence of our experiments is greatly improved by using Conditional Value at Risk (CVaR) as the accumulation function, instead of a more traditional energy minimisation. In this setting, only the top-performing samples are taken into consideration when evaluating the merit of the QAOA parameters. CVaR only cares about “high value” samples with low energy, and high energy outliers are discarded. CVaR makes intuitive sense since the output of QAOA is a single sample from the final distribution rather than any average-case answer. Recent studies have justified CVaR both analytically [30] and empirically [30, 31, 32].
Finally, we find that we can noticeably increase the speed of simulations by improving the efficiency of the classical sample processing. A naive sample-by-sample evaluation of the incurred cost takes time comparable to the circuit simulation time. Using bitwise operations at the byte level reduces the classical processing time to a negligible fraction of the quantum simulation time.
@@ Acknowledgements
The authors would like to thank the entire QPG team for their valuable support and discussions throughout the process of this work. Work on the Quantum Pangenomics project was supported by Wellcome Leap as part of the Q4Bio Program. JB was additionally funded by the Wellcome Trust [Grant number 220540/Z/20/A]. SS was also supported by the Royal Society University Research Fellowship.
@@
[…]
@@
[…]
@@
We implemented a local graph preprocessing step that observes cases where a node has multiple input and output edges and at least one input and one output has non-zero edge weight and non-zero node weight. In this case, any additional edges with zero weights can be deemed to be superfluous and removing them will not break graph connectivity. The results can be seen in  table 4, which demonstrates a significant drop in the number of breaks usually and a slight increase in percentage covered and used.
@@
Table 4: Results of the MQLib solver running for 30 seconds on 51 graphs aligning 5 sequences each, with and without edge trimming enabled. Results sequence with consensus optimisation.
[…]

### v2

@@ 4.7 Mapping QUBO problems to quantum optimisation
The circuit is constructed from a layer of Hadamard gates followed by alternating cost and mixer layers. The cost layer is $U_{C}(\gamma)=\exp(-i\gamma H)$. There are many choices for a suitable mixer layer; we use $U_{M}(\beta)=\exp(-i\beta\sum_{i}X_{i})$. The number of such layers is a hyperparameter $p$. In the limit as $p\rightarrow\infty$, QAOA approaches adiabatic quantum computing, which is known to be universal for quantum computing [1]. In contrast, for $p=1$, QAOA is known to be classically simulable under certain circumstances [15]. In the absence of noise, increasing $p$ should increase the convergence rate of QAOA. In practice, the effects of noise mean that a moderate $p$ value is most likely to be effective.
@@ 4.8 Classical simulation of QAOA
We find that the convergence of our experiments is greatly improved by using Conditional Value at Risk (CVaR) as the accumulation function, instead of a more traditional energy minimisation. In this setting, only the top-performing samples are taken into consideration when evaluating the merit of the QAOA parameters. CVaR only cares about “high value” samples with low energy, and high energy outliers are discarded. Recent studies have justified CVaR both analytically [4] and empirically [4, 20, 7].
Key points
  • Using pangenomes to resolve short-read data avoids linear reference bias that is unavoidable in traditional reference-guided mapping.
  • Compared to de novo assembly, pangenome-guided assemblies report far fewer contigs and cover a greater amount of the true genome, with only a small increase in the number of false joins.
  • Our novel optimisation-based approach achieves comparable results to bleeding-edge exhaustive methods, and will remain feasible as problem size increases.
  • Our methods are designed to run efficiently on quantum computers as quantum technology improves; already, quantum annealing methods achieve good results on moderate-sized instances.
@@ Acknowledgements
The authors would like to thank the entire QPG team for their valuable support and discussions throughout the process of this work.
Funding
Work on the Quantum Pangenomics project was supported by Wellcome Leap as part of the Q4Bio Program. JC was supported by the EPSRC. JB was additionally funded by the Wellcome Trust [Grant number 220540/Z/20/A]. SS was also supported by the Royal Society University Research Fellowship.
@@
Biographical note
Josh Cudby is a PhD student at the University of Cambridge and a Research Assistant at the University of Oxford. His research focuses on quantum algorithms for bioinformatics.
@@
S2 Future research directions
One drawback of the current optimisation formulation is the lack of edge information. When we have graphs with multiple paths through the same set of nodes, it can be impossible to distinguish nodes A+ B+ C+ from A+ B- C+. Pathfinder combines edge weights with node weights to resolve these conflicts. Experiments using pathfinder with and without edge weights have shown that using this information has a minor impact on most statistics, but can reduce the number of false joins by up to a third.
While it is straightforward to include an edge weight term in the abstract optimisation cost functions of eqs. 1, 2 and 3, it is not clear how to translate such terms into a QUBO. The information of whether an edge was traversed at a certain time is inherently quadratic, given by terms $x_{t,\,i}x_{t+1,\,j}$. To obtain a cost function that is minimised when the number of traversals of an edge is close to the weight of that edge requires terms like $\sum_{t}\big(x_{t,\,i}x_{t+1,\,j}-w(i,\,j)\big)^{2}$ for each $(i,\,j)\in E$. These higher-order terms incur overheads in the number of variables if they are re-expressed as quadratic terms. A local graph preprocessing step to remove zero-weight edges that do not break graph connectivity is likely to achieve an approximate equivalent without the need to incorporate edge data directly into the QUBO model.
All graph path finding methods are impacted by the quality of the input kmer counts. Further improvements can be made to these tools, particularly concerning non-unique mapping, which would improve the quality of all solvers. However, there is further missing information yet to be considered, where larger sequence fragments trace a path through many small nodes, forming longer-range constraints rather than the nearest neighbour ones encoded in the graph edge counts.
The kmer mapping software used to generate node weights does not use Illumina read-pairing metrics. In a traditional linear reference alignment, the read pairing is an important factor for resolving some repeats, and the same could also be applied to pangenome mapping. Pairing information could potentially also be used in inferring paths, utilising knowledge of the template size distribution.
One possible direction that improves resource usage and also allows edge weight data to be included is the use of Higher-order Unconstrained Binary Optimisation (HUBO) variants. HUBO formulations are more expressive than QUBO, with terms containing products of arbitrarily many binary variables; the downside is that (classically) solving them is generally more difficult and less well-explored in the literature. Notably, however, HUBO may not have such a solution overhead when mapped to QAOA algorithms on quantum computers. In that setting, each layer of the cost Hamiltonian now contains multi-qubit $Z$ rotations rather than single- and two-qubit rotations only. These gates need to be compiled down to 2-qubit gates, but constructions that are linear in the number of qubits involved exist [13]. In this setting, the inclusion of edge weight data requires the implementation of more multi-qubit interactions, which in turn increases the circuit depth. Notably, there are no overheads in the number of qubits, which is a major limiting factor in current-term quantum hardware. We could also consider Prog-QAOA [2] formulations, which directly encode optimisation problems into QAOA cost Hamiltonians, without going through a QUBO or HUBO step. We expect that either of these 2 encodings could achieve $\mathcal{O}(N\log(N))$ qubit scaling, at the cost of increased circuit depth.
$C_{G}(W)=\sum_{v\in V}\log(L(v))\Bigl(\#W(v)-w(v)\Bigr)^{2},$ (7)
where $L(v)$ is the length of the sequence stored in node $v$.
S3 Details of node weight assignment methods
@@
A schematic representation of the kmer2node annotation strategy on a small graph.
Figure S1: A pictorial view of kmer2node processing on a graph built from 4 short example genomes. On the top left we have a fully complete pangenome capturing variation at the individual base level (minigraph-cactus style). Top right is a lossy graph with small variants collapsed into representative nodes (minigraph style). Bottom left is produced by aligning the full length genomes used to build the graph back to the graph and tracking which sequence substrings align to each node. Bottom right is after recursively walking back through the graph from each node to generate k1k-1 prefix bases.
The pangenome graphs we process come from minigraph, which does not produce new nodes for minor sequence variation, leading to imprecise mapping. This is beneficial as it reduces the complexity of the graphs to resolve, making the problem more amenable to the quantum environment, but it also means kmer matching may miss some hits. We do, however, have the original input genomes that were used to build the pangenome. By aligning these back to the pangenome graph with GraphAligner, we can get a path of node names along with a CIGAR string. Combined, these yield the exact sequence from the training genomes that covers each node rather than the single one reported by minigraph. This ancillary “nodeseq” file is also indexed, improving the mapping accuracy of kmer2node. See fig. S1 for a pictorial demonstration of these steps.
@@
We implemented a local graph preprocessing step that observes cases where a node has multiple input and output edges and at least one input and one output has non-zero edge weight and non-zero node weight. In this case, any additional edges with zero weights can be deemed to be superfluous and removing them will not break graph connectivity. The results can be seen in  table 4, which demonstrates a significant drop in the number of breaks and a slight increase in percentage covered and used.
@@
Table 4: Results of the MQLib solver running for 30 seconds on 51 graphs aligning 5 sequences each, with and without edge trimming enabled. Results are with the consensus optimisation applied.
S5 Preliminary results of hybrid assembly and kmer mapping